OpenStudy (kitkat1):

Medal for best Answer!!!! Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? (1 point) f(n) = 28(0.5)n f(n) = 56(0.5)n− 1 f(n) = 56(0.5)n f(n) = 112(0.5)n− 1

OpenStudy (anonymous):

You want to know what is the relationship between the n values and the f function

OpenStudy (kitkat1):

I think its either the B or C

OpenStudy (anonymous):

By looking at the answers, since the is multiples of 0.5, it tells you you have to divide.

OpenStudy (anonymous):

0.5 is just 1/2

OpenStudy (kitkat1):

Okay whats the differnece between B and C?

OpenStudy (kitkat1):

So whats the answer?

OpenStudy (anonymous):

Well you have to divide each by 2

OpenStudy (anonymous):

So 56/2=?

OpenStudy (anonymous):

28!

OpenStudy (anonymous):

28/2=? 14!

OpenStudy (kitkat1):

28

OpenStudy (anonymous):

Since each divides evenly by 1/2

OpenStudy (kitkat1):

Are u sure?

OpenStudy (anonymous):

so what i just told you tells you that each one is 0.5

OpenStudy (anonymous):

but the initial value is 56

OpenStudy (anonymous):

You have to test the points.

OpenStudy (anonymous):

if you plug in n=1 into B it gives you the correct answer. Sorry.

OpenStudy (anonymous):

Its B

OpenStudy (kitkat1):

Okaycan u help me with one more?

OpenStudy (anonymous):

Yesh

OpenStudy (kitkat1):

Jake has proved that a function, f(x), is a geometric sequence. How did he prove that? (1 point) He showed that an explicit formula could be created. He showed that a recursive formula could be created. He showed that f(n) ÷ f(n − 1) was a constant ratio. He showed that f(n) − f(n − 1) was a constant difference. Im pretty sure its the 3rd one but just want to make sure

OpenStudy (kitkat1):

Am I right its the 3rd one?

OpenStudy (anonymous):

an = ar(n – 1)

OpenStudy (kitkat1):

3rd one?

OpenStudy (anonymous):

Yes it is the third

OpenStudy (anonymous):

since ar=ar^2 so ar^2/ar=r

OpenStudy (kitkat1):

Okay I have one more question that can u check if I got it right?

OpenStudy (anonymous):

constant ratio of r

OpenStudy (anonymous):

yesh

OpenStudy (kitkat1):

A store had 250 bottles of water. Each week, 40% of the bottles were sold and 48 new bottles arrived in shipments. Which recursive function best represents the number of bottles of water in the store, given that f(0) = 250? f(n) = f(n − 1) ⋅ 0.6 + 48, n > 0 f(n) = 250 − f(n − 1) ⋅ 0.4 + 48, n > 0 f(n) = f(n − 1) ⋅ 0.4 + 48, n > 0 f(n) = 250 − f(n − 1) ⋅ 0.6 + 48, n > 0 My answer is D.

OpenStudy (anonymous):

Yes D makes sense

OpenStudy (kitkat1):

Okay thank you

OpenStudy (anonymous):

But why would it not be B

OpenStudy (anonymous):

.6 would suggest that 60% of the bottles are being sold.